You have a laser and want to estimate the wavelength. You decide to use a reflective diffraction grating to do this. The light is incident on the grating at 20 degrees with respect to normal. Using a grating with a groove distance of 5 mm, the first-order diffraction, is found at an angle of-12.5 degrees. Determine the wavelength of the source.
Answer #1

We can use Grating equation to calculate wavelength of light source. m\lambda = d (sin\thetai+ sin\thetar) . Where m is order of diffraction, \thetai is incident angle and \thetar is diffracted angle.

m =1, \thetai = 20°, \thetar = -12.5° and d =5 mm = 5×10-3m

\lambda = 5×10-3m ( sin20° + sin(-12.5°)) = 5×10-3×0.125581m

\lambda = 0.6279×10-3m = 6.279×10-4 m . (answer)